# Molecular Identification and Peak Assignment: Leveraging Multi-Level Multimodal Alignment on NMR

### v1

@@ 3.1.5 Baselines
In contrast to communicative mechanism of K-M3AID, one of the baselines is established without the utilization of a communication channel. Additionally, we explore two alternative baselines where the knowledge-guided instance-wise discrimination in the node-level alignment module is replaced with either strong-pair-based (SP-ID) or weak-pair-based (WP-ID) instance-wise discrimination. SP-ID enforces a precise match in node-level (atom-peak) alignment, ensuring that only correct pairs established during the training process are considered. On the other hand, WP-ID broadens the matching criteria, allowing for multiple matches within a specified threshold set for the distance of their corresponding parts per million (ppm, referenced in Section  6.4). The mathematical definitions of strong pair and weak pair are as follows:
@@
Here, $A_{i}\in A$ represents atoms, $P_{j}\in P$ represents peaks, and $th$ is the abbreviation for the threshold. Additionally, $ppm_{i}$ denotes the parts per million for the $i^{th}$ Carbon Atom, and $ppm_{j}$ represents the corresponding parts per million for the $j^{th}$ peak.
@@ 3.2.1 Validation Performance
The K-M3AID model demonstrates an impressive validation accuracy of 95.5% for molecule-spectrum alignments within the graph-level alignment module. In direct comparison, K-M3AID surpasses alternative models such as SP-ID, WP-ID, and the model without a communication mechanism, exhibiting superior performance by a margin ranging from about 1% to 6% in the graph-level alignment module (refer to Table 1). Notably, SP-ID outperforms WP-ID significantly, and as the matching criteria threshold widens, the performance of the latter degrades.
In the context of peak-atom alignment within the node-level alignment module, K-M3AID and the model without a communication mechanism exhibit comparable accuracies (refer to Table 1). However, K-M3AID showcases slightly better stability across 5-fold cross-validation. Moreover, K-M3AID demonstrates superiority in peak-atom alignment when compared to SP-ID and WP-ID. This superiority may arise from the inherent limitations of both strong and weak pair definitions, which is failing to precisely calibrate the diverse relationships among the elements. This finding is further supported by the significant decreases in the accuracy of peak-atom alignment as the threshold of weak pair increases.
Table 1: Batch-wise validation accuracy (%) of K-M3AID and baselines with $epochs=200$.
@@ 3.2.2 Performance on Zero-Shot Molecular Retrieval
We conduct a systematic evaluation of the effectiveness of our K-M3AID model, comparing it with baseline models in the zero-shot molecular retrieval task across datasets of varying magnitudes. Detailed results are presented in Table 2. The K-M3AID model consistently attains an impressive top-1 accuracy of approximately 95.8% in molecular retrieval when the molecular reference library comprises 100 entries. This performance surpasses that of alternative mechanisms such as SP-ID (95.3%), WP-ID (92.9%), and No Communication (94.8%). As the molecular reference library expands to 1000 entries, the K-M3AID model exhibits notable superiority, achieving accuracy levels of 80.4%, 1.8%, 8.7%, and 2.8% higher than SP-ID, WP-ID, and No Communication mechanisms, respectively. The advantage of K-M3AID becomes even more pronounced when the library size reaches 10,000 entries. In this scenario, K-M3AID yields 46.3% at top-1 accuracy, showcasing advancements of 10.5%, 13.6%, and 6.2% over SP-ID, WP-ID, No Communication mechanisms, respectively. Even with larger molecular reference libraries, such as 100,000 and 1,000,000 entries, K-M3AID consistently outshines SP-ID, WP-ID, and No Communication mechanisms. These compelling results distinguish the K-M3AID model as an exceptional choice in scenarios demanding robust performance in molecular retrieval tasks.
Table 2: Zero-shot molecular retrieval at top 1(%) accuracy across datasets of varying sizes. (For more statistics, such as top 5(%), top 10(%) and top 25(%), please refer to Appendix Table  C.1.)
@@ 3.2.4 Performance on Zero-Shot Peak Assignment
K-M3AID model demonstrates a validation accuracy surpassing 90% for peak assignment (peak-atom alignment) within the node-level alignment module after 200 epochs (see Figure 2.A). Notably, the model achieves a 100% accuracy rate in 74.1% of molecules containing fewer than 10 carbon atoms (see Figure 2.B). For molecules with carbon atom counts ranging from 10 to 20, the model attains 100% accuracy in 37.2% of cases (see Figure 2.C). Furthermore, it achieves an accuracy exceeding 80% in more than 50% of cases pertaining to molecules containing more than 20 carbon atoms (see Figure 2.D). Additionally, to further illustrate the power of K-M3AID on peak assignment, we present two complex natural product molecules featuring multiple rings (4 and 4, respectively) and stereogenic (chiral) centers (6 and 8, respectively) in Figure E.1.
@@
K-M3AID demonstrates superior performance in peak assignment compared to SP-ID and WP-ID. Our case studies reveal that the limitations of SP-ID and WP-ID become particularly pronounced in two scenarios: 1) when local contexts of specific atoms exhibit a high degree of similarity; 2) when certain atoms display symmetric mapping within the same molecule. In the former scenario, exemplified by molecular A in Figure  3, atom 0 and atom 4 are secondary carbons (attaching to 2 carbons and 2 hydrogens), nearly symmetric on the same 5-member ring, corresponding to the peak position measured in ppm of 27.0 and 29.8, respectively. (For the definition of ppm, please refer to Section 6.4.) The similar local content of these two atoms fools SP-ID and WP-ID. Meanwhile, atom 1 and atom 3 are tertiary carbons (attaching to 3 carbons and 1 hydrogen), nearly symmetric on the same 5-member ring, corresponding to the peak position measured in ppm of 54.5 and 44.1, respectively. Only WP-ID fails to distinguish and align them. In the later scenario, exemplified by molecular B in Figure  3, there exist instances one-to-one and one-to-many for atomic-level alignment within the molecular configuration. Both SP-ID and WP-ID method misaligns certain atoms with other atoms with small ppm differences (less than 3 ppm in this case), rather than aligning them with themselves or their symmetric counterparts. In contrast, K-M3AID approach excels in both scenario by discerning each one of the atoms, which is attributed to the full utilization of ppm difference distance learning. (For additional cases, please refer to Appendix Figure E.2)
Figure 3: Case study of peak assignment. Yellow cells in PPM difference represent the ground truth alignment, and red cross represents the wrong alignment. For the definition of ppm, please refer to 6.4. For additional cases, please refer to Appendix Figure E.2
@@ 4 Related Work
Multimodal Instance-Wise Discrimination: As mentioned in the prelimtaries, instance discrimination (Le-Khac et al. (2020a); Zolfaghari et al. (2021); Morgado et al. (2021); Liu et al. (2023)), an important part of contrastive learning, distinguishes individual instances without explicit class labels. Moving into multimodal contrastive learning, it can be categorized into two general approaches: strong-pair-based (van den Oord et al. (2019); Jaiswal et al. (2021); Liu et al. (2023)) and weak-pair-based (Salakhutdinov & Hinton (2007); Frosst et al. (2019); Liang et al. (2021)) instance-wise discrimination. The strong-pair-based such as NCE method enforces a precise one-to-one correspondence for real samples with artificially generated noise samples. An example of positive pair can be a noise-added picture of zebra with the text description of zebra. Instead of one-to-one correspondences, weak-pair-based approach relaxes the positive pairs to more boarder semantic correspondences. An example of positive pair can be a picture of zebra with the text description of horse, but not with the text description of tiger.
Multimodal Meta-Alignment: Within the scope of multimodal alignment, multimodal meta-alignment is a novel method to align representation spaces using paired cross-modal data with different similarity while ensuring quick generalization to new tasks across different modalities (Liang et al. (2021)). This approach can be observed at different levels, including the intermediate and fundamental (irreducible) element level. Examples of this method at the intermediate level can be seen in research on Cross-Modal Generalization (Chen et al., 2017; Li et al., 2020; Liang et al., 2021; Zhang et al., 2021) and Livestreaming Product Recognition (Yang, 2023). While these studies showcase how multimodal meta-alignment operates at a broad objective level, the application of multimodal meta-alignment at the most fundamental element level remains underexplored in current research.
@@ 5 Conclusion and Future Work
In this paper, we introduced K-M3AID (Knowledge-Guided Multi-Level Multimodal Alignment with Instance-Wise Discrimination) framework, incorporating graph-level alignment and node-level alignment. Its effectiveness was demonstrated through multiple zero-shot tasks: molecular retrieval, isomer recognition and peak assignment. The significance of knowledge-guided instance-wise discrimination is underscored through various metrics and case studies. Moreover, the findings from molecular retrieval and isomer recognition underscored the favorable influence of the node-level alignment on the graph-level alignment. It emphasizes the successful integration of meta-learning within our hierarchical alignment framework. While our framework achieves an atomic-level alignment overall accuracy of 100% for 55% of cases, it drops significantly to 9.8% when handling molecules with more than 20 carbon atoms. Presently, our graph encoder operates on 2D molecular graphs with basic node and edge features. This implementation potentially constrains its ability to generate precise node embeddings for distinguishing atoms in highly complex scenarios. Future developments could benefit from incorporating a 3D-based graph, holding substantial potential to enhance performance in such complex situations.
6 Our Method
In this section, we firstly present the architecture of K-M3AID framework, an end-to-end system designed for multi-level multimodal alignment. Then, we introduce the constrastive learning loss in K-M3AID along with the principles of Knowledge-Guided Instance-Wise Discrimination.
6.1 Architecture
The K-M3AID framework is a dual-CLIP architecture (see Figure 1), which consists of three critical components: graph-level alignment module, node-level alignment module and communication channel. The graph-level alignment module adapts a gradient-asymmetric CLIP mechanism. While two unimodal encoders work in conjunction, only the from-scratch graph encoder (GIN, Xu et al. (2018)) undergoes dynamic training throughout the process, the pre-trained spectrum encoder (Yang et al. (2021)) remains fixed. Both encoders are complemented by dedicated projection layers, which facilitate the mapping of embeddings into a joint space. The node-level alignment module adapts a gradient-symmetric CLIP mechanism. It is equipped with two from-scratch unimodal encoders, node encoder and peak encoder, as well as their dedicated projection layers. The graph encoder in graph-level alignment module shares part of the weights with node encoder in node-level alignment module, serving as the communication channel.
6.2 Contrastive Learning Loss
The synergy between these two modules is pivotal, collectively contributing to our loss function, expressed as
$L=CL_{graph}+CL_{node},$ (3)
$=-\text{log}(\text{softmax}(\delta(x_{i},y_{i}))$ (5)
Thus, the total $CL_{graph}$ is expressed as following:
$CL_{graph}=\frac{1}{N}\displaystyle\sum_{1\leq i\leq N}CL_{graph}(i)$ (6)
Figure 4: Knowledge span mechanism. $K_{i}$ and $K_{j}$ represents the corresponding knowledge span label for $i^{th}$ and $j^{th}$ items.
Knowledge Span, in which we define as some continuous and domain-specific feature exhibiting natural ordering and offering guidance, can potentially offer valuable insights into the contrastive learning process. As such, we propose a novel and general approach to contrastive learning called Knowledge-Guided Instance-Wise Discrimination (see Figure  4). This approach expands the scope of contrastive learning from confined comparisons (pre-determined negative and positive pairs) to unrestricted comparisons (no need for pre-determination). This extension removes the necessity of explicitly defining such pairs, thus mitigating the potential introduction of human bias.
6.4 Chosen Knowledge Span-ppm
${}^{13}$ NMR uncovers molecular structures by providing the chemical environments of carbon atoms and their magnetic responses to external fields, and quantifies these features in parts per million (ppm) relative to a reference compound like tetramethylsilane (TMS), simplifying comparisons across experiments. Thus, continuous peak positions, measured in ppm, can serve as a robust knowledge span to facilitate instance-wise discrimination for this contrastive learning task.
For node-level alignment module, $\mathcal{A}$ is the set of node embeddings for Carbon atoms in molecular graph modality and $\mathcal{B}$ is the set of peak embeddings for respective Carbon atoms in NMR modality, $\mathcal{K}$ is the set of ppm value for each corresponding Carbon atom in $\mathcal{A}$ and $\mathcal{B}$. Suppose $ppm_{i}$ is the ppm for the $i^{th}$ Carbon Atom and $ppm_{j}$ is the corresponding ppm for $j^{th}$ peak. $d(\cdot,\cdot)$ is then defined as following:
$d(\mathcal{K}_{i},\mathcal{K}_{j})=d(ppm_{i},ppm_{j})$ $=softmax(\frac{\tau_{2}}{|ppm_{i}-ppm_{j}|+\tau_{1}})$ (14)
where $\tau_{1}$ and $\tau_{2}$ are temperature hyper-parameter. For further discussion of selection about $\tau_{1}$ and $\tau_{2}$, please refer to Appendix B.2. Then, the final form of contrastive loss for node-level alignment according to Equation 11 and Equation 13 is as following:
$CL_{node}=-\frac{1}{|\mathcal{M}|}\sum_{\begin{subarray}{c}1\leq j\leq|\mathcal{M}|\end{subarray}}d(ppm_{i},ppm_{j})\log\frac{e^{d_{E}(\mathcal{A}_{i},\mathcal{B}_{j})}}{\sum_{\begin{subarray}{c}1\leq k\leq|\mathcal{M}|\end{subarray}}e^{d_{E}(\mathcal{A}_{i},\mathcal{B}_{k})}}$ (15)
@@ Appendix A Revisiting Knowledge Span Guided Loss
Appendix B Further ablation study about parameter choices
B.1 Ablation study about the choice of GIN structure and projection.
We choose GIN(Xu et al. (2018)) as our graph encoder. By Table B.1, ”GIN Depth” signifies the number of layers in the GIN, ”GIN Embedding Dim” denotes the dimensionality of the embeddings generated by the GIN model, and ”Projection Dim” indicates the resulting dimensionality after transforming the GIN-produced embeddings. In particular, the best performance is observed when the GIN model has 5 layers, GIN Embedding Dim is 128, and projection Dim is 512.
Table B.1: GIN structure and projection ablation study

### v2

@@ 4.2 Baselines
Strong-Pair-based Instance-Wise Discrimination: We explore an alternative baseline where the knowledge-guided instance-wise discrimination in the node-level alignment module is replaced with strong-pair-based instance-wise discrimination (denoted as SP-ID). SP-ID enforces a precise match in node-level (atom-peak) alignment, ensuring that only correct pairs established during the training process are considered. The mathematical definition of a strong pair is as follows:
@@
where $i$, $j$ represent the indices of nodes (atoms).
Weak-Pair-based Instance-Wise Discrimination: We replace the knowledge-guided instance-wise discrimination with weak-pair-based instance-wise discrimination (denoted as WP-ID) in the node-level alignment module. WP-ID broadens the matching criteria of SP-ID, allowing for multiple matches within a specified threshold set for the distance of their corresponding parts per million (ppm, referenced in Section  4.1). The mathematical definition of a weak pair is as follows:
@@
where $i$, $j$ represent the indices of nodes (atoms).
@@ 4.3.1 Validation Performance
The K-M3AID model showcases an impressive validation accuracy of 95.5% in aligning molecules with spectra within the graph-level alignment module. In direct comparison, K-M3AID outperforms alternative models such as SP-ID, WP-ID, and the model without a communication mechanism (No Comm.), demonstrating superior performance with a margin ranging from approximately 1% to 6% in the graph-level alignment module (refer to Table 1). Notably, SP-ID significantly outperforms WP-ID, and as the matching criteria threshold widens, the performance of the latter deteriorates.
In the context of peak-atom alignment within the node-level alignment module, K-M3AID and the model without a communication mechanism exhibit comparable accuracies (refer to Table 1). However, K-M3AID showcases slightly better stability across 5-fold cross-validation. Moreover, K-M3AID demonstrates superiority in peak-atom alignment when compared to SP-ID and WP-ID. This superiority may arise from the inherent limitations of both strong and weak pair definitions, which fail to precisely calibrate the diverse relationships among the elements. This finding is further supported by the significant decreases in the accuracy of peak-atom alignment as the threshold of weak pair increases.
Table 1: Batch-wise validation accuracy (%) of K-M3AID and baselines with $epochs=200$. For WP-ID, the threshold is configured at 1, 5, and 10 ppm.
@@ 4.3.2 Performance on Zero-Shot Molecular Retrieval
We conduct a systematic evaluation of the effectiveness of our K-M3AID model, comparing it with baseline models in the zero-shot molecular retrieval task across datasets of varying magnitudes. Detailed results are presented in Table LABEL:tab:pubmed-retrieval. The K-M3AID model consistently attains an impressive top-1 accuracy of approximately 95.8% in molecular retrieval when the molecular reference library comprises 100 entries. This performance surpasses that of alternative mechanisms such as SP-ID (95.3%), WP-ID (92.9%), and No Comm. (94.8%). As the molecular reference library expands to 1000 entries, the K-M3AID model exhibits notable superiority, achieving accuracy levels of 80.4%, 1.8%, 8.7%, and 2.8% higher than SP-ID, WP-ID, and No Comm. mechanisms, respectively. The advantage of K-M3AID becomes even more pronounced when the library size reaches 10,000 entries. In this scenario, K-M3AID yields 46.3% at top-1 accuracy, showcasing advancements of 10.5%, 13.6%, and 6.2% over SP-ID, WP-ID, and No Comm. mechanisms, respectively. Even with larger molecular reference libraries, such as 100,000 and 1,000,000 entries, K-M3AID consistently outshines SP-ID, WP-ID, and No Comm. mechanisms. These compelling results distinguish the K-M3AID model as an exceptional choice in scenarios demanding robust performance in molecular retrieval tasks.
Table 2: Zero-shot molecular retrieval at top 1 accuracy across datasets of varying sizes. (For more statistics, such as top 5, top 10 and top 25, please refer to Appendix Table  D.1. For similarity comparison between the molecules and the Top 1 neighbor by different retrieval methods, please consult Appendix Table  D.2.)
@@ 4.3.4 Performance on Zero-Shot Peak Assignment
The K-M3AID model demonstrates a validation accuracy surpassing 90% for peak assignment (peak-atom alignment) within the node-level alignment module after 200 epochs (see Figure 3.A). Notably, the model achieves a 100% accuracy rate in 74.1% of molecules containing fewer than 10 carbon atoms (see Figure 3.B). For molecules with carbon atom counts ranging from 10 to 20, the model attains 100% accuracy in 37.2% of cases (see Figure 3.C). Furthermore, it achieves an accuracy exceeding 80% in more than 50% of cases pertaining to molecules containing more than 20 carbon atoms (see Figure 3.D). Additionally, to further illustrate the power of K-M3AID on peak assignment, we present two complex natural product molecules featuring multiple rings (4 and 4, respectively) and stereogenic (chiral) centers (6 and 8, respectively) in Figure F.1.
@@
K-M3AID demonstrates superior performance in peak assignment compared to SP-ID and WP-ID. Our case studies reveal that the limitations of SP-ID and WP-ID become particularly pronounced in two scenarios: 1) When local contexts of specific atoms exhibit a high degree of similarity. 2) When certain atoms display symmetric mapping within the same molecule.
In the former scenario, exemplified by molecular A in Figure 4, atom 0 and atom 4 are secondary carbons (attaching to 2 carbons and 2 hydrogens), nearly symmetric on the same 5-member ring, corresponding to the peak position measured in ppm of 27.0 and 29.8, respectively (for the definition of ppm, please refer to Section 4.1). The similar local content of these two atoms fools SP-ID and WP-ID. Meanwhile, atom 1 and atom 3 are tertiary carbons (attaching to 3 carbons and 1 hydrogen), nearly symmetric on the same 5-member ring, corresponding to the peak position measured in ppm of 54.5 and 44.1, respectively. Only WP-ID fails to distinguish and align them.
In the latter scenario, exemplified by molecular B in Figure 4, there exist instances one-to-one and one-to-many for atomic-level alignment within the molecular configuration. Both SP-ID and WP-ID methods misalign certain atoms with other atoms with small ppm differences (less than 3 ppm in this case), rather than aligning them with themselves or their symmetric counterparts. In contrast, the K-M3AID approach excels in both scenarios by discerning each one of the atoms, which is attributed to the full utilization of ppm difference distance learning. (For additional cases, please refer to Appendix Figure F.2)
Figure 4: Case study of peak assignment. Yellow cells in PPM difference represent the ground truth alignment, and red cross represents the wrong alignment. For the definition of ppm, please refer to 4.1. For additional cases, please refer to Appendix Figure F.2
@@ 5 Related Work
Multimodal Instance-Wise Discrimination: As mentioned in the preliminaries, instance discrimination (Le-Khac et al., 2020a; Zolfaghari et al., 2021; Morgado et al., 2021; Liu et al., 2023), an important part of contrastive learning, distinguishes individual instances without explicit class labels. Transitioning into multimodal contrastive learning, it can be categorized into two general approaches: strong-pair-based (van den Oord et al., 2019; Jaiswal et al., 2021; Liu et al., 2023) and weak-pair-based (Salakhutdinov & Hinton, 2007; Frosst et al., 2019; Liang et al., 2021) instance-wise discrimination. The strong-pair-based approach, such as the Noise Contrastive Estimation (NCE) method, enforces a precise one-to-one correspondence for real samples with artificially generated noise samples. An example of a positive pair can be a noise-added picture of a zebra with the text description of a zebra. Instead of one-to-one correspondences, the weak-pair-based approach relaxes the positive pairs to broader semantic correspondences. An example of a positive pair can be a picture of a zebra with the text description of a horse but not with the text description of a tiger.
Multimodal Meta-Alignment: Within the realm of multimodal alignment, multimodal meta-alignment is a novel method for aligning representation spaces using paired cross-modal data with different similarity levels while ensuring quick generalization to new tasks across different modalities (Liang et al., 2021). This approach can be observed at different levels, including the intermediate and fundamental (irreducible) element level. Examples of this method at the intermediate level can be found in research on Cross-Modal Generalization (Chen et al., 2017; Li et al., 2020; Liang et al., 2021; Zhang et al., 2021) and Livestreaming Product Recognition (Yang et al., 2023). While these studies showcase how multimodal meta-alignment operates at a broad objective level, the application of multimodal meta-alignment at the most fundamental element level remains underexplored in current research.
@@ 6 Conclusion and Future Work
In this paper, we introduced the K-M3AID (Knowledge-Guided Multi-Level Multimodal Alignment with Instance-Wise Discrimination) framework, incorporating both graph-level and node-level alignment. Its effectiveness was demonstrated through multiple zero-shot tasks, including molecular retrieval, isomer recognition, and peak assignment. The significance of knowledge-guided instance-wise discrimination is underscored through various metrics and case studies. Moreover, the findings from molecular retrieval and isomer recognition highlight the favorable influence of node-level alignment on graph-level alignment. This emphasizes the successful integration of meta-learning within our hierarchical alignment framework. While our framework achieves an atomic-level alignment overall accuracy of 100% for 55% of cases, it drops significantly to 9.8% when handling molecules with more than 20 carbon atoms. Currently, our graph encoder operates on 2D molecular graphs with basic node and edge features. This implementation potentially constrains its ability to generate precise node embeddings for distinguishing atoms in highly complex scenarios. Future developments could benefit from incorporating a 3D-based graph, holding substantial potential to enhance performance in such complex situations.
Accessibility
The code and dataset will be made available upon the date of publication.
@@ Appendix A Revisiting Knowledge Span Guided Loss
Appendix B Experimental Setting
B.1 Pre-training
Dataset: We use ${}^{13}$ NMR spectra of about 20,000 molecules sourced from nmrshiftdb2 (Steinbeck et al., 2003)), a public access database that contains NMR spectra of organic molecules. In the collected dataset, molecule are aligned with their respective ${}^{13}$ NMR spectra, and atomic alignments with peaks are also included. Notably, the dataset contains 12,771 molecules with fewer than 10 carbon atoms, 7,043 molecules featuring carbon atom counts ranging from 10 to 20, and 1,138 molecules incorporating more than 20 carbon atoms. The quality of the dataset was further validated by experienced organic chemists. We randomly sample 80% of the molecules for training and the rest for evaluation.
B.2 Zero-Shot Molecular Retrieval
Dataset: We randomly collected 1 million molecules from PubChem (Kim et al., 2023) to form a molecular reference library. Subsequently, we carefully selected 1000 spectra, ensuring that they had not appeared in the training dataset, from an external dataset to serve as query spectra. Following this, the corresponding molecules associated with these 1000 spectra were added into the existing reference library.
Evaluation: We perform molecular retrieval using each of the selected spectra to determine if the correct corresponding molecular entity can be retrieved from the reference library. The model’s performance is assessed at top-1 accuracy, as well as at accuracy of top 5, top 10, and top 25.
B.3 Zero-Shot Isomer Recognition
Dataset: We categorize isomers from the validation dataset to guarantee their absence from the training dataset. To assess effectiveness, we perform isomer recognition on each isomer group containing at least 10 molecules. Within the same group, isomers may be structural or spatial isomers of each other. Structural isomers refer to molecules with the same molecular formula but different structural arrangements of atoms, resulting in distinct chemical structures. On the other hand, spatial isomers, also known as stereoisomers, have the same molecular formula and arrangement of atoms but differ in the spatial orientation of their atoms in three-dimensional space, leading to different stereoisomeric forms. (For an elucidation of an isomer group and in-depth insights into isomers with NMR, please refer to the details provided in Appendix E.)
B.4 Zero-Shot Peak Assignment
Appendix C Further ablation study about parameter choices
C.1 Ablation study about the choice of GIN structure and projection.
We choose GIN(Xu et al., 2018) as our graph encoder. By Table C.1, ”GIN Depth” signifies the number of layers in the GIN, ”GIN Embedding Dim” denotes the dimensionality of the embeddings generated by the GIN model, and ”Projection Dim” indicates the resulting dimensionality after transforming the GIN-produced embeddings. In particular, the best performance is observed when the GIN model has 5 layers, GIN Embedding Dim is 128, and projection Dim is 512.
Table C.1: GIN structure and projection ablation study